# Inference for Regression with

### v2

@@ 5.2 CEO Behavior
The estimation sample consists of 916 CEOs, each of whom participated in a survey that recorded features of time use in each 15-minute interval of a given week, e.g. Monday 8am-8:15am, Monday 8:15am-8:30am, and so forth. The recorded categories are (1) the type of activity (meeting, public event, etc.); (2) duration of activity (15m, 30m, etc.); (3) whether the activity is planned or unplanned; (4) the number of participants in the activity; (5) the functions of the participants in the activity (HR, finance, suppliers, etc.). In total there are 654 unique combinations of these categories observed in the data. We let $x_{i,j}$ denote the number of times feature $j$ appears in the time use diary of CEO $i$. The average value of $C_{i}$ is 88.4, with a minimum of 2 and a maximum of 222. Bandiera et al., (2020) uses LDA with $K=2$ dimensions to organize the time use data. The authors refer to the separate distributions over time use combinations $\beta_{1}$ and $\beta_{2}$ as pure behaviors. The share of CEO $i$’s time devoted to pure behavior $1$, $\theta_{i}$, is referred to as the CEO index.
The authors use the following inference procedure. First, estimate LDA on the time use data using the collapsed Gibbs sampler of Griffiths and Steyvers, (2004), then form an estimate $\hat{\theta}_{i}$ based on the posterior means. They then use $\hat{\theta}_{i}$ as an input into productivity regressions where $Y_{i}$ is the log of firm $i$ sales, and $\textbf{q}_{i}$ is a vector of firm observables. Further, they separately analyzed which CEO and firm characteristics are associated with behaviors by regressing $\hat{\theta}_{i}$ on a vector of characteristics $\textbf{g}_{i}$.
@@
As demonstrated both theoretically and through the simulation exercise, the key quantity that governs the relative importance of sampling error and measurement error is $\kappa$. In the context of the CEO behavior data, the empirical analog of $\kappa$ is the product of the square root of the number of observations (CEOs) and the average value of the inverse of the number of activities per CEO. This value is 0.44, which is close to the lowest value of $\kappa$ in the simulation exercise. This suggests that the two-step approach should perform relatively well in this application. To further test our theory, we also estimate the model using data where we first sample 10% of the activities for each CEO, without replacement, with the minimum of one. This scenario could represent a researcher observing only half of a workday for each CEO, instead of a full five-day workweek. Such sampling increases the analogue of $\kappa$ to 4.26, which is near the middle value of $\kappa$ in the simulation exercise, indicating that we should expect the two-step approach to perform poorly under these conditions.
@@ 6 Conclusion
The leading strategy for analyzing unstructured data uses two steps. First, quantitative representations of unstructured data are extracted in an information retrieval step. Second, the derived quantitative representations are plugged into downstream econometric models, with the representations treated as regular numerical data for the purposes of estimation and inference. This paper highlights, both theoretically and empirically, a previously unrecognized problem with this popular two-step strategy: measurement error introduced in the first step leads to biased estimates and invalid inference for downstream regression coefficients. The degree of bias, and therefore the degree to which it distorts inference, depends on the relative importance of measurement error and sampling error, but it can be material in applications. To guard against it, we propose a robust inference method based on maximum likelihood estimation of the information retrieval and regression models jointly. We implement this one-step strategy using Hamiltonian Monte Carlo deployed on modern hardware. This strategy outperforms the two-step strategy in simulations and generates quantitatively important differences in a leading application.
While we develop theoretical arguments within a simple regression model, we posit the same effects will arise in more elaborate downstream econometric models. For example, an emerging line of research uses text-derived sentiment indices as inputs into forecasting models with a vector autoregressive or dynamic factor structure. Straightforward extensions of our theoretical arguments can be used to show how error in the indices will bias coefficient estimates and limit the effectiveness of these forecasting methods. More constructively, the one-step strategy can be used to enhance the performance of these forecasting methods. Likewise, the industrial organization literature is increasingly using embedded representations of firms and products to characterize market behavior and demand with structural models. Our one-step strategy can be used to mitigate bias introduced by measurement error in the embeddings in these and other cases where the downstream model has a likelihood formulation. Going forward, it is important to establish for which specific information retrieval methods and econometric models does measurement error most severely affect inference.
@@
Finally, we note there are limits on the scalability of Hamiltonian Monte Carlo, even when fully optimized. When one confronts a vast amount of data, alternative approaches for approximating the joint distribution in the one-step strategy must be used. One popular choice in computer science is variational inference (VI; Jordan et al., 1998, Wainwright and Jordan, 2008) which has recently seen applications in economics (Bonhomme, 2021, Olenski and Sacher, 2022, Mele and Zhu, 2023). [^19] Recently, scalable extensions of HMC that are based on stochastic approximations of gradients have shown initial promise (Dang et al., 2019). From an implementation perspective, VI is no more complicated than HMC because it too can be formulated within probabilistic programming languages that rely on automatic differentiation (Hoffman et al., 2013). [^20] For an example, see https://num.pyro.ai/en/stable/tutorials/tbip.html which illustrates a VI implementation of the Text-Based Ideal Points model (Vafa et al., 2020). However, VI has fewer statistical guarantees than HMC. In ongoing work, we are studying how to best perform scalable inference in the one-step strategy with massive data.
@@ A.1 Proofs for Section 2
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\theta}_{i}^{k_{1}}\theta_{i}^{k_{2}}-\mathbb{E}[\hat{\theta}_{i}^{k_{1}}\theta_{i}^{k_{2}}]\right|>t\right)\leq\frac{\mathbb{E}[\hat{\theta}_{i}^{2k_{1}}\theta_{i}^{2k_{2}}]}{t^{2}n}\leq\frac{1}{t^{2}n}\,.$
Consider the denominator term in $T_{1,n}$ and $T_{2,n}$. By Chebyshev’s inequality, we have
@@
$\mathbb{E}\left[\sqrt{n}(\hat{\theta}_{i}-\theta_{i})(\hat{\theta}_{i}-\mathbb{E}[\theta_{i}])\right]$ $=\mathbb{E}\left[\sqrt{n}(\hat{\theta}_{i}-\theta_{i})^{2}\right]$
@@
A second application of Chebyshev’s inequality gives
@@
because $(\hat{\bar{\theta}}-\mathbb{E}[\theta_{i}])\times\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\to_{p}0$. Note that $\mathbb{E}[\varepsilon_{i}(\hat{\theta}_{i}-\theta_{i})]=0$ because $Y_{i}$ and $(X_{i},C_{i})$ are independent conditional on $\theta_{i}$ and both $\varepsilon_{i}$ and $\hat{\theta}_{i}-\theta_{i}$ have conditional (on $\theta_{i}$) mean zero. Hence by Chebyshev’s inequality, for any $t>0$ we have
@@
$\hat{Y}_{i}-Y_{i}=\frac{\hat{\theta}_{i}-\theta_{i}}{\theta_{i}(1-\theta_{i})}+\frac{(2\theta_{i}-1)(\hat{\theta}_{i}-\theta_{i})^{2}}{2\theta_{i}^{2}(1-\theta_{i})^{2}}+\frac{(-3\tilde{\theta}_{i}^{2}+3\tilde{\theta}_{i}-1)(\hat{\theta}_{i}-\theta_{i})^{3}}{3\tilde{\theta}_{i}^{3}(1-\tilde{\theta}_{i})^{3}}$
@@
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{(-3\tilde{\theta}_{i}^{2}+3\tilde{\theta}_{i}-1)(\hat{\theta}_{i}-\theta_{i})^{3}}{3\tilde{\theta}_{i}^{3}(1-\tilde{\theta}_{i})^{3}}(Z_{i}-\mathbb{E}[Z_{i}])\right|\\ \leq D\max_{1\leq i\leq n}|\hat{\theta}_{i}-\theta_{i}|\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{\theta}_{i}-\theta_{i})^{2}|Z_{i}-\mathbb{E}[Z_{i}]|\right)$
@@ A.2 Proofs for Section 3
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]=\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right),$
@@
$\mathbb{E}\left[\left.(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}\right|C_{i},\bm{\theta}_{i}\right]=\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right).$
@@
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]$ $=\frac{1}{C_{i}^{2}}\mathbb{E}\left[\left.\mathbf{x}_{i}\!\!\!\;\mathbf{x}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]$
@@
$=\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right),$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]-\mathbb{E}\left[\frac{1}{C_{i}}\right]\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\right\|\to_{p}0.$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right\|\to_{p}0$
where $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}$ exists with probability approaching one by Assumption 1(i)-(ii). Each element of $\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}$ is bounded between 00 and $1$, so we may deduce by Chebyshev’s inequality that
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right\|\to_{p}0.$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\to_{p}\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right],\quad\quad\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\Bigg{)}^{-1}\to_{p}\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]^{-1}.$
@@
The first result follows by the law of large numbers (LLN). The second result then follows because the rank condition on $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ implies $\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]$ has full rank. ∎
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right].$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right)\right\|\\ \leq\left(\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|\right)\times\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{q}_{i}\|\to_{p}0.$
Moreover, $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\to_{p}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}$ by Assumption 1(i)-(ii). Let $\mathbf{X}_{i}=\mathbf{q}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}[\mathbf{q}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}]$ and let $D$ denote the dimension of $\mathbf{q}_{i}$. Then
@@
by Assumption 1(iv) and the fact that $\hat{\mathbf{p}}_{i}$ is in the simplex. Hence, $\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}\to_{p}0$. The result follows by Slutsky’s theorem, noting that $\mathbb{E}[\mathbf{q}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}]=\mathbb{E}[\mathbf{q}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\mathbf{B}$. ∎

### v3

@@ 5.2 CEO Behavior
The estimation sample consists of 916 CEOs, each of whom participated in a survey that recorded features of time use in each 15-minute interval of a given week, e.g. Monday 8am-8:15am, Monday 8:15am-8:30am, and so forth. The recorded categories are (1) the type of activity (meeting, public event, etc.); (2) duration of activity (15m, 30m, etc.); (3) whether the activity is planned or unplanned; (4) the number of participants in the activity; (5) the functions of the participants in the activity (HR, finance, suppliers, etc.). In total there are 654 unique combinations of these categories observed in the data. We let $x_{i,j}$ denote the number of times feature $j$ appears in the time use diary of CEO $i$. The average value of $C_{i}$ is 88.4, with a minimum of 2 and a maximum of 222. Bandiera et al., (2020) uses LDA with $K=2$ dimensions to organize the time use data. The authors refer to the separate distributions over time use combinations $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ as pure behaviors. The share of CEO $i$’s time devoted to pure behavior $1$, $\theta_{i}$, is referred to as the CEO index.
The authors use the following inference procedure. First, estimate LDA on the time use data using the collapsed Gibbs sampler of Griffiths and Steyvers, (2004), then form an estimate $\hat{\theta}_{i}$ based on the posterior means. They then use $\hat{\theta}_{i}$ as an input into productivity regressions where $Y_{i}$ is the log of firm $i$ sales, and $\textbf{q}_{i}$ is a vector of firm observables. Further, they separately analyze which CEO and firm characteristics are associated with behaviors by regressing $\hat{\theta}_{i}$ on a vector of characteristics $\textbf{g}_{i}$.
@@
As demonstrated both theoretically and through the simulation exercise, the key quantity that governs the relative importance of sampling error and measurement error is $\kappa$. In the context of the CEO behavior data, the empirical analog of $\kappa$ is the product of the square root of the number of observations (CEOs) and the average value of the inverse of the number of activities per CEO. This value is 0.44, which is close to the lowest value of $\kappa$ in the simulation exercise. This suggests that the two-step approach should perform relatively well in this application. To further test our theory, we also estimate the model using data where we first sample 10% of the activities for each CEO, without replacement, with a minimum of one. This scenario could represent a researcher observing only half of a workday for each CEO, instead of a full five-day workweek. Such sampling increases the analogue of $\kappa$ to 4.26, which is near the middle value of $\kappa$ in the simulation exercise, indicating that we should expect the two-step approach to perform poorly under these conditions.
@@
Note: Each point represents the mean posterior estimate of a single CEO’s index, $\hat{\theta}_{i}$. The blue line is the local polynomial fit (with confidence intervals) obtained with ‘ggplots’s’ ‘geom_smooth’ with default parameters.
@@ 6 Conclusion
The leading strategy for analyzing unstructured data uses two steps. First, quantitative representations of unstructured data are extracted in an upstream information retrieval step. Second, the derived quantitative representations are plugged into downstream econometric models, with the representations treated as regular numerical data for the purposes of estimation and inference. This paper highlights, both theoretically and empirically, a previously unrecognized problem with this popular two-step strategy: measurement error introduced in the first step leads to biased estimates and invalid inference for downstream regression coefficients. The degree of bias, and therefore the degree to which it distorts inference, depends on the relative importance of measurement error and sampling error, but it can be material in applications. To guard against it, we propose a robust inference method based on joint maximum likelihood estimation of the IR and regression models. Joint estimation is straightforward using HMC and modern probabilistic programming languages. This strategy outperforms the two-step strategy in simulations and generates quantitatively important differences in a leading application.
Implementing the one-step strategy requires formulating a likelihood. Latest-generation machine-learning- and AI-based approaches to information retrieval increasingly use neural networks with no obvious structure that yields a likelihood. While implementing the one-step strategy may not be possible in these settings, the measurement error problem does not thereby disappear: it simply becomes harder to characterize. Such approaches are often given statistical foundations following their adoption and, as this process plays out, the scope for the one-step strategy will expand accordingly. [^17] One illustrative example is the popular word2vec model for producing word embeddings. The original model (Mikolov et al., 2013b , Mikolov et al., 2013a ) had no statistical interpretation but yielded word representations that nevertheless captured semantic relationships well. Word2vec has subsequently been adopted by economists as part of the two-step strategy, for example to measure occupation-level exposure to technological change (Kogan et al., 2019) and emotionality in political speech (Gennaro and Ash, 2022). In parallel, a literature has developed likelihood-based interpretations of embeddings (Arora et al., 2016, Dieng et al., 2020, Ruiz et al., 2020) which could in principle be adapted for use in the one-step strategy.
@@
Finally, we note there are limits on the scalability of HMC, even when fully optimized. When one confronts a vast amount of data, alternative approaches for approximating the joint distribution in the one-step strategy must be used. One popular choice in computer science is variational inference (VI; Jordan et al., 1998, Wainwright and Jordan, 2008) which has recently seen applications in economics (Bonhomme, 2021, Olenski and Sacher, 2022, Mele and Zhu, 2023). VI is no more complicated to implement because it too can be formulated within probabilistic programming languages that rely on automatic differentiation (Hoffman et al., 2013). However, VI has fewer statistical guarantees than HMC because it uses an approximate likelihood in place of the true likelihood. In ongoing work, we are studying how to best perform scalable inference in the one-step strategy with massive data.
@@ A.1 Proofs for Section 2
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\theta}_{i}^{k_{1}}\theta_{i}^{k_{2}}-\mathbb{E}[\hat{\theta}_{i}^{k_{1}}\theta_{i}^{k_{2}}]\right|>t\right)\leq\frac{\mathbb{E}[\hat{\theta}_{i}^{2k_{1}}\theta_{i}^{2k_{2}}]}{t^{2}n}\leq\frac{1}{t^{2}n}\,.$ (16)
Consider the denominator term in $T_{1,n}$ and $T_{2,n}$. By inequality (16), we have
@@
$\mathbb{E}\left[\sqrt{n}(\hat{\theta}_{i}-\theta_{i})(\hat{\theta}_{i}-\mathbb{E}[\hat{\theta}_{i}])\right]$ $=\mathbb{E}\left[\sqrt{n}(\hat{\theta}_{i}-\theta_{i})^{2}\right]$
@@
A second application of inequality (16) gives
@@
because $(\bar{\hat{\theta}}-\mathbb{E}[\theta_{i}])\times\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\to_{p}0$. Note that $\mathbb{E}[\varepsilon_{i}(\hat{\theta}_{i}-\theta_{i})]=0$ because $Y_{i}$ and $(X_{i},C_{i})$ are independent conditional on $\theta_{i}$ and both $\varepsilon_{i}$ and $\hat{\theta}_{i}-\theta_{i}$ have conditional (on $\theta_{i}$) mean zero. Hence by Chebyshev’s inequality, for any $t>0$ we have
@@
$\hat{Y}_{i}-Y_{i}=\frac{\hat{\theta}_{i}-\theta_{i}}{\theta_{i}(1-\theta_{i})}+\frac{(2\theta_{i}-1)(\hat{\theta}_{i}-\theta_{i})^{2}}{2\theta_{i}^{2}(1-\theta_{i})^{2}}+\frac{(3\tilde{\theta}_{i}^{2}-3\tilde{\theta}_{i}+1)(\hat{\theta}_{i}-\theta_{i})^{3}}{3\tilde{\theta}_{i}^{3}(1-\tilde{\theta}_{i})^{3}}$
@@
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{(3\tilde{\theta}_{i}^{2}-3\tilde{\theta}_{i}+1)(\hat{\theta}_{i}-\theta_{i})^{3}}{3\tilde{\theta}_{i}^{3}(1-\tilde{\theta}_{i})^{3}}(Z_{i}-\mathbb{E}[Z_{i}])\right|\\ \leq D\max_{1\leq i\leq n}|\hat{\theta}_{i}-\theta_{i}|\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{\theta}_{i}-\theta_{i})^{2}|Z_{i}-\mathbb{E}[Z_{i}]|\right)$
@@ A.2 Proofs for Section 3
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]=\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right),$
@@
$\mathbb{E}\left[\left.(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}\right|C_{i},\bm{\theta}_{i}\right]=\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right).$
@@
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]$ $=\frac{1}{C_{i}^{2}}\mathbb{E}\left[\left.\mathbf{x}_{i}^{\phantom{T}}\!\!\!\;\mathbf{x}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]$
@@
$=\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right),$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]-\mathbb{E}\left[\frac{1}{C_{i}}\right]\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\right\|\to_{p}0.$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right\|\to_{p}0$
where $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}$ exists with probability approaching one by Assumption 1(i)-(ii). Each element of $\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}$ is bounded between 00 and $1$, so we may deduce by Chebyshev’s inequality that
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right\|\to_{p}0.$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathbb{E}\left[\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\to_{p}\mathbb{E}\left[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right],\quad\quad\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\Bigg{)}^{-1}\to_{p}\mathbb{E}\left[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right]^{-1}.$
@@
The first result follows by the law of large numbers (LLN). The second result then follows because the rank condition on $\mathbb{E}\left[\bm{\xi}_{i}^{\phantom{T}}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ implies $\mathbb{E}\left[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right]$ has full rank. ∎
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\theta}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right].$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}^{\phantom{T}}\!\!\!\;\mathbf{q}_{i}^{T}\right)\right\|\\ \leq\left(\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|\right)\times\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{q}_{i}\|\to_{p}0.$
Moreover, $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\to_{p}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}$ by Assumption 1(i)-(ii). Let $\mathbf{X}_{i}=\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}]$ and let $D$ denote the dimension of $\mathbf{q}_{i}$. Then
@@
by Assumption 1(iv) and the fact that $\hat{\mathbf{p}}_{i}$ is in the simplex. Hence, $\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}\to_{p}0$. The result follows by Slutsky’s theorem, noting that $\mathbb{E}[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}]=\mathbb{E}[\mathbf{q}_{i}^{\phantom{T}}\!\!\!\;\bm{\theta}_{i}^{T}]\mathbf{B}$. ∎